A daredevil is shot out of a cannon at 45.0° to the horizontal with an initial speed of 28.0 m/s. A net is positioned a horizontal distance of 51.0 m from the cannon. At what height above the cannon should the net be placed in order to catch the daredevil?
Answer #1

Solution:

This question is based on the concept of:

1. Projectile motion

2. Height attained by the projectile

The detailed solution is explained below:

Find time of the projectile motion
x = Vx(t)
51 = 28cos45(t)
t = 2.575 seconds

Find height (y) of the projectile:
y = Vy(t) + (1/2)(a)(t^2)
y = 28(sin45)(2.575) - (4.905)(2.575)^2
y = 18.46 m <-- Answer (Height above the cannon)

Thanks!!!